A sample solution containing sulfate ions (SO42−) was analyzed and was found to contain 2.5 mol of sulfate ions (SO42−).
Calculate the equivalent of sulfate ions (SO42−) present in the soluti
Answer #1

Ans.:- 5.0 equivalent.

Since, 1 mole of SO4^2- = 2 moles of charge = 2 equivalent of SO4^2-.

so, 1 mole of SO4^2- = 2 equivalent of SO4^2-.

Therefore, 2.5 mole of SO4^2- = (2.5 mole of SO4^2-) * (2 equivalent SO4?^2-)/ ( 1mole of SO4^2-)

= 5.0 equivalent of So4^2- Ans